What area under the standard normal curve falls outside the Z values -2.5 and 2.5?

*the answer is not 0.9876 as someone else already incorrectly answered. please give the correct answer and show your work :) thank you!
Answer #1

Solution :

Given that,  

Using standard normal table ,

= 1 - P(-2.5 < z < 2.5)

= 1 - P(z < 2.5) - P(z < -2.5)

= 1 - P(0.9938 - 0.0062)

= 1 - 0.9876

= 0.0124

Area = 0.0124